OpenStudy (anonymous):

A body hangs from a spring balance supported from the roof of an elavator. If the elavator has an upward acceleration of 3ms−2 and the balance reads 50 N, what is the true weight of the body?

OpenStudy (amistre64):

isnt it just calculated using newtons second law?

OpenStudy (amistre64):

F = ma m = F/a W = mg = Fg/a